Three charges are placed on the corners of an equilateral triangle (10μC at the top, -10μC at the lower right, and 5μC at the lower left). Find the magnitude and the direction of the electric field at the center (Point P) of the triangle. The length of one of the sides is 0.5 m.

(a) 28.8 x 106 N/C, θ = 36o

(b) 1.93 x 106 N/C, θ = -44o

(c) 1.02 x 107 N/C, θ = -57o

(d) zero

I'd really appreciate the help!
Answer #1

43 30 ら

r = distance of each corner from the center = L/(2 sqrt(3)) = 0.5 /(2 sqrt(3)) = 0.14 m

Ea = k qa/r2 = (9 x 109) (10 x 10-6)/(0.14)2 = 4.6 x 106 N/c

Eb = k qb/r2 = (9 x 109) (5 x 10-6)/(0.14)2 = 2.3 x 106 N/c

Ec = k qc/r2 = (9 x 109) (10 x 10-6)/(0.14)2 = 4.6 x 106 N/c

Net electric field along the X-direction is given as

Ex = Ec Cos30 + Eb Cos30 = (4.6 x 106 ) Cos30 + (2.3 x 106 ) cos30 = 5.98 x 106 N/C

Net electric field along the Y-direction is given as

Ey = - Ea - Ec Cos60 + Eb Sin60 = - (4.6 x 106 ) - (4.6 x 106 ) Cos60 + (2.3 x 106 ) sin30 = - 5.75 x 106 N/C

net electric field is given as

E = sqrt(Ex2 + Ey2) = sqrt((5.98 x 106 )2 + ( - 5.75 x 106 )2) = 8.3 x 106 N/C

direction :

tan-1(Ey/Ex) = tan-1(5.75 x 106 /(5.98 x 106 )) = - 44